Ford claims that the 2016 Mustang GT provides 425 HP (horsepower) of power. A random sample of 17 Mustang GTs was taken and the horse power ratings were followed an approximately normal model with a mean of 407.2 HP and a standard deviation of 41.96 HP. Find the critical value for a 99% confidence interval.
Answer #1

Solution :

n = 17

Degrees of freedom = df = n - 1 = 17 - 1 = 16

At 99% confidence level the t is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2,df = t0.005,16 = 2.921

Critical value = 2.921